25 men can unload a truck in 2 hours.how many hours should it take 15 men to unload the truck
a)4.33 b)3.33 c)5.33 d)7.33

Answer by Edwin McCravy(20054)

25 men can unload a truck in 2 hours.how many hours should it take 15 men to unload the truck
25 men can unload a truck in 2 hours

5 men can unload a truch in 10 hours

15 men can unload a truck in 10/3 hours = 3 1/3 hours

WORKERS*TIME     WORKERS*TIME
------------  =  ------------
   JOBS             JOBS

  25*2   15*T
  ---- = -----
    1      1

    50 = 15T

 50/15 = T

  10/3 = T

 3 1/3 = T

  3.33 = T 

Edwin